y=4x-3. Find these 3 points on this line. (1, ),( ,)),(-4, )
Answer by graphmatics(170)

The line is given by the equation y=4x-3. For (1,) the x is 1 so the corresponding y is y=4*1-3=1. This point is (1,1). For the (4,) we have y=4*4-3=13 so the point (4,13) is on the line. For (,) it looks like we just select an x between 1 and 4, say x=2. So for x=2 y=4*2-3=5 so the point (2,5) is on the line.